Q18 (1 point). The table below shows the demand and supply schedules for peanuts. Suppose the government imposes a 12 cent taQ19 (1 point). The table below shows the demand and supply schedules for peanuts. Suppose the government imposes a 12 cent tanot sure from my answers ^^
Answer #1

18 ans is B. Total there is a tax of 60 cents. So PB-Ps=0.6 so new quantity=60 because at q=60, Pb=4.8 and Ps=4.2

19. Tax Revenue=t*Q=0.6(60)=36